1) Cu immersed in a solution 0,001 M CuSO4. Determine Er of Cu-electrode (using the Nernst equation) and then calculate the EMF of a voltaic cell, versus the saturated calomel electrode (SCE) with E=0,24V

Write electrochemical equations for the oxidation and reduction processes occurring during the corrosion of Al in air-free sulphuric acid (H2SO4). Write the anode and cathode reactions.
1
Expert's answer
2019-10-17T07:32:30-0400

E=Eo + 0.059/n x lg(ox/red)= 0.16 + 0.059/2 x lg 0.001= 0.0715V.

EMF= 0.24 - 0.0715 = 0.1685 .